OpenStudy (anonymous):

how do you prepare 425 ml of 12.5% NaCl

OpenStudy (anonymous):

We can assume that the density of the solution is the same as that of water - 1 g/ml. So, we have \[m(sol) = d*V = 1*425 = 425 g\] So, we know that 12,5 % of that solution is NaCl, which means: \[ m(NaCl) = w*m(sol) = 0.125 * 425 = 53.125 g \] Which means the remaining mass of the solution is due to water: \[ m(H2O) = m(sol) - m(NaCl) = 371.875 g \]